The table below is of hourly labor productivity (i.e., output per employee).

Rice

Wheat

Labor Productivity: China

10 kg/hour

4 kg/hour

  

Labor Productivity: Japan

8 kg/hour

6 kg/hour

  1. What are the absolute advantages in this example?
  2. What are the comparative advantages in this example? How do you know this?
  3. After a trade liberalization what would China export and import? Why?
  4. After a trade liberalization what would Japan gain? Why?
Answer #1

a.

Absolute advantage can be defined as the ability of a Nation to produce a greater quantity of a goods or services compare to its rivals by using the same amount of resources.

Hence China has absolute advantage in Rice because it needs less resources to produce same output compare to another country. Here china is able to produce more quantity of output with the same resources compare to Japan.

Hence Japan has absolute advantage in Wheat because it needs less resources to produce same output compare to another country. Here Japan is able to produce more quantity of output with the same resources compare to china.

b.

The opportunity costs can be defined as the lost units of output of other goods for producing an additional unit of output of a good.

A country has comparative advantage in the production of that good in which it has lower opportunity cost.

Opportunity cost of Rice in China=4/10

=0.4 Kg of wheat

Opportunity cost of wheat in China=10/4

=2.5 Kg of rice

Opportunity cost of Rice in Japan=6/8

=0.75 Kg of wheat

Opportunity cost of wheat in Japan=8/6

=1.33 Kg of rice

Since opportunity cost of Rice is lower in china compare to Japan, so China has comparative advantage in Rice production.

Similarly Since opportunity cost of wheat is lower in Japan compare to china, so Japan has comparative advantage in wheat production.

c.

After a trade liberalization China would export Rice and import wheat because it has comparative advantage in rice.

d.

After a trade liberalization Japan would import Rice and export wheat because it has comparative advantage in wheat.